# A review on combinatorial approach to aggregation

### v1

@@ II.3 Probability distribution over state space
The main aim of the combinatorial analysis is to derive the probability distribution over the state space of the system, i.e., the probability $P(\mathrm{\Omega})$ of finding the coagulating system in a given state $\mathrm{\Omega}$,
$P(\mathrm{\Omega})=\frac{W(\mathrm{\Omega})}{Z},$ (7)
@@
where $W(\mathrm{\Omega})$ is the thermodynamic probability and $Z=\sum_{\mathrm{\Omega}}{W(\mathrm{\Omega})}$ stands for the normalizing factor Fronczak_2018 . Allowed states of the system are not equiprobable and $P(\mathrm{\Omega})\neq\mathrm{const}$, which is a result of non–equilibrium character of the process.
Eq. (7) will be later used to derive an average number of clusters of given size $s$, $\langle n_{s}\rangle=\sum_{\Omega}{n_{s}(\Omega)P(\Omega)}$, and its standard deviation.
@@
Throughout the following sections, we will derive expressions to calculate the thermodynamic probabilities $W(\Omega)$ and the corresponding $Z$ factor. We will take advantage of the three observations: (i) the set of monomers can be divided into subsets in a particular number of ways, (ii) a cluster of a given size $g$ can be created in a particular number of ways, $x_{g}$ (the number of possible histories of a cluster), and (iii) coagulation acts for a particular cluster can be distributed in different time steps.
II.4 Remark on subscripts $s$ and $g$
Here, we must comment on subscripts denoting the size of a cluster. Up to this point, naturally, we have been using $s$ to denote the size; thus, $\langle n_{s}\rangle$ is the average number of clusters of size $s$. However, in the following derivations, we will write $x_{g}$ to denote the number of ways to create a cluster of size $g$. This switching, strange at first glance, will become justified much later in the text when a final expression for $\langle n_{s}\rangle$ will arise (cf. Eq. (37)). Then, the subscript $s$ will stand for the size of interest (a particular size for which we find the number of clusters), and $g$, used in the argument of Bell polynomials, will cover the range from $1$ to Nk+1N-k+1 (i.e., to $t+1$) being dependent on the largest-cluster size that could grow in the system.
Using the same subscript would make distinguishing between these two (in general different) numbers difficult.
@@ II.4 Remark on subscripts $s$ and $g$
Figure 4: a) A graphical representation of the cluster selection process in the case of coagulation, in which the probability of two particles merging does not depend on their sizes. Our task is to create a particle of size $g$, i.e., consisting of $g$ monomers. The monomers are distinguishable (numbered for clarity). In the first step, we combine two available $g$ particles (here $g=4$). In the next step, we combine two further freely selected clusters. We merge the particles until we obtain a single cluster. In this way, we implement one of $x_{4}=4!3!/2^{3}=18$ ways to create a cluster of size $4$. b) Diagram illustrating all possible growth histories of such a particle of size $4$. The number of different growth histories, $x_{g}$, equals the number of different paths drawn by arrows and leading through different states of the diagram. c) A tree corresponding to the bold path in the diagram in b). Single particle growth history may be illustrated as a rooted tree, with leaf nodes standing for monomeric units, internal nodes representing the history-dependent transition states of the cluster, and the root node being the final step in the growth process. Each vertical line represents a state of the system in the particular time step. Illustration by the authors.
II.5 Distributing monomers into clusters
@@
In this section, we will calculate the number of ways to distribute monomers into the given number of clusters (i.e., partitioning the system of $N$ monomers into $k$ subsets) Fronczak_2018 . Let $N$ distinguishable objects be divided into $k$ non-empty and disjoint subsets of $c_{i}>0$ elements each, where $\sum^{k}_{i=1}{c_{i}=N}$. As we can choose $c_{1}$ in $\binom{N}{c_{1}}$ ways, $c_{2}$ in $\binom{N-c_{1}}{c_{2}}$ ways, there are
@@
of such partitions; $n_{g}\geq 0$ stands for the number of subsets of size $g$, with the largest subset size equal to Nk+1N-k+1. The change of the multiplication index $i$ to $g$ means that we no longer multiply over the subsets (clusters) but over the sets of subsets (clusters) of a given size determined by $g$.
For example, let us consider a system of $N=5$ monomers grouped into $k=2$ clusters of sizes 2 and 3. Remember that all monomers are distinguishable (i.e., we can assign numbers from $1$ to $N=5$ to these particles). Possible grouping methods are presented in Fig. 3. There are ten ways of doing so.
@@
$N!\prod^{N-k+1}_{g=1}{\frac{1}{n_{g}!}{\left(\frac{x_{g}}{g!}\right)}^{n_{g}}}.$ (10)
Summing Eq. (10) over all integers $\left\{n_{g}\right\}$ specified by Eq. (6) one obtains the partial Bell polynomial $B_{N,k}\left(\left\{x_{g}\right\}\right)$, an important fact that will be used in Sec. II.9. The Reader is encouraged to consult Appendix A for the definition and interpretation of partial Bell polynomials.
Internal state $x_{g}$ of the cluster corresponds to the number of ways in which the cluster of size $g$ could be created in the system (in other words, the number of possible histories of the cluster). In general, $x_{g}$ varies for different kernels. The following subsection shows how to obtain $x_{g}$ for any given kernel.
To sum up this part, Eq. (10) gives the number of possible ways of partitioning $N$ monomers into $k$ clusters for a given state $\mathrm{\Omega}$ of the system. However, it implicitly assumes that all the clusters in the system arose at once, which is not valid, as only one merging act is allowed per one time step. Coagulation acts corresponding to different clusters may alternate, increasing the number of ways in which a given (final) microstate can be created. This issue leads to the need to calculate the number of possible distributions of merging acts in time (see Sec. II.7).
II.6 Calculating number $x_{g}$ of internal states of cluster of given size for constant kernel
@@
where $x_{h}$ and $x_{g-h}$ represent the numbers of ways to create two clusters of size $h$ and (gh)(g-h) that joined and became a cluster of size $g$.
@@
The second binomial factor, $\binom{g-2}{h-1}$, covers the fact that coagulation acts resulting in clusters $h$ and (gh)(g-h) could occupy different possible time steps (cf. Fig. 5).
The sum in Eq. (12) is taken over all possible pairs of clusters that can join to become a cluster of size $g$. The factor of 1/2 is used to prevent double counting of the coagulation acts. This factor is arbitrary in the sense of the final result, $\langle n_{s}\rangle$, which would not change if the factor of 1/2 was removed. More comments on this factor may be found in Ref. Lepek_2019 .
In Sec. II.8, we will show how to obtain Eq. (11) from Eq. (12).
II.7 Distributing joining acts in time
Figure 5: A graphical representation of the evolution of a coagulation system consisting of $N=7$ monomers, in which two particles with masses 3 and 4 were formed. There are ten possible sequences of coagulation acts leading to the forming of such two clusters. To create them, we need five acts of coagulation, i.e., five time steps. The smaller cluster, size 3, is created in two time steps, and the larger cluster, size 4, is created in three time steps. In part (a), we can see that the black dots, representing the coagulation acts that lead to the formation of a cluster of size 3, and the white dots, representing the coagulation acts that lead to the formation of a cluster of size 4, can be ordered in ten different ways in time. Part (b) of the figure shows two of the ten possible system evolution scenarios. Coagulation acts associated with forming a cluster of size $h$ can occur in h1h-1 time steps out of the total number of N2=5N-2=5 steps needed to form clusters of sizes 3 and 4. Reprinted from Lepek_2021 , with permission from Elsevier.
As signaled in the previous subsections, a given state can arise as a result of different sequences of intermixed coagulation acts corresponding to different clusters. So far, we have not considered the fact that coagulation acts leading to the creation of a cluster can occur in different steps in time. Therefore, the distribution of coagulation acts in time is the second origin of combinatorial equations.
@@ II.7 Distributing joining acts in time
$\begin{split}\left[\binom{t}{1}\binom{t-1}{1}\dots\binom{t-n_{2}+1}{1}\right]\dots\left[\binom{t-n_{2}}{2}\binom{t-n_{2}-2}{2}\dots\binom{t-n_{2}-2(n_{3}-1)}{2}\right]\dots{\\ }=\frac{t!}{{\left(1!\right)}^{n_{2}}{\left(2!\right)}^{n_{3}}\dots\left(\left(g-1\right)!\right)^{n_{g}}\dots}=t!\prod^{N-k+1}_{g=1}{\frac{1}{{\left(\left(g-1\right)!\right)}^{n_{g}}}}.\end{split}$ (13)
For a better understanding of Eq. (13), consider, as before, the system of $N=5$ monomers after $t=3$ time steps. Let us assume that $k=2$ clusters exist in that system: one cluster of size 2 (dimer) and one cluster of size 3 (trimer). From Eq. (13) we have
$t!\prod^{N-k+1}_{g=1}{\frac{1}{{\left(\left(g-1\right)!\right)}^{n_{g}}}}=\frac{3!}{{\left(0!\right)}^{0}{\left(1!\right)}^{1}{\left(2!\right)}^{1}{\left(3!\right)}^{0}}=3,$ (14)
@@
$\frac{7!}{{\left(0!\right)}^{0}{\left(1!\right)}^{0}{\left(2!\right)}^{1}{\left(3!\right)}^{1}{\left(4!\right)}^{0}{\left(5!\right)}^{0}}=10.$ (15)
@@
Sequences that can be observed for that system are illustrated in Fig. 5.

### v2

@@ II.3 Probability distribution over state space
The main aim of the combinatorial analysis is to derive the probability distribution over the state space of the system, i.e., the probability $P(\Omega,t)$ of finding the coagulating system in a given state $\mathrm{\Omega}(t)$, cf. Eq. (3),
$P(\Omega,t)=\frac{W(\Omega,t)}{Z(t)},$ (8)
@@
where $W(\Omega,t)$ is the thermodynamic probability (a number of microstates that correspond to a single macrostate) of a (macro)state $\Omega(t)$ and $Z(t)=\sum_{\mathrm{\Omega}}{W(\Omega,t)}$ stands for the normalizing factor analogous to partition function in statistical mechanics Fronczak_2018 . Allowed states of the system are not equiprobable and $P(\mathrm{\Omega})\neq\mathrm{const}$, which is a result of non-equilibrium character of the process.
An important comment must be given here. Eq. (8), written analogously as for the microcanonical distribution, is only as assumption that will later allow do derive further quantities: an average number of clusters of given size $s$, $\langle n_{s}\rangle=\sum_{\Omega}{n_{s}(\Omega)P(\Omega)}$, and its standard deviation, and probability distribution $P(\Omega,t)$ itself. Eq. (8) assumes that the probability of a given state is proportional to the number of microstates realizing such a given state. Thus, Eq. (8) describes a system where all the ways to create a state $\Omega$ (possible realizations of the system over time), their number being $W(\Omega,t)$, are equally probable. In other words, by the equiprobability of those ways, we mean that if there are multiple ways (microstates) of ending up at the same state $\Omega$ (macrostate), then all the routes have equal probability of occurring.
The assumption of equiprobable ways of obtaining the state $\Omega$ is entirely valid for the constant kernel, for which it was initially used in Ref. Fronczak_2018 . However, it does not hold for arbitrary kernels, being only an approximation for them. Further comments on that issue may be found in Sec. V.1 and in Appendix C.
In Eq. (8), we formally shall use notation $P(\Omega(t))$, $W(\Omega(t))$, but as it would be obfuscating in reading, we used notation with a comma. The sense of the notation is that, at time $t$, there is a set of possible states $\Omega$ that could exist in the system. In further writing, we will usually omit explicit dependence on $t$ as the state $\Omega$, although formally dependent on $t$, constitutes a more fundamental and unambiguous quantity, being a base for further analysis. We believe it will also shorten the expressions.
@@
Throughout the following sections, we will derive expressions to calculate the thermodynamic probabilities $W(\Omega,t)$ and the corresponding $Z(t)$ factor. We will take advantage of the three observations: (i) the set of monomers can be divided into subsets in a particular number of ways, (ii) a cluster of a given size $g$ can be created in a particular number of ways, $x_{g}$ (the number of possible histories of a cluster), and (iii) coagulation acts for a particular cluster can be distributed in different time steps.
Figure 3: a) Possible ways to group a set of $N=5$ elements into two clusters of sizes 2 and 3. All particles are distinguishable. There are ten ways to do so (cf. Eq. (9)). b) Illustration of the change of the multiplication index from $i$ to $g$, cf. Eq. (9), meaning that we no longer enumerate over the subsets (i.e., over clusters which are assigned with their respective numbers $i=1,2,3,4,5$) but over the sets of subsets (i.e., the sets of clusters) of a given size determined by $g$ (for this example, $g=1,2,3$, $\Omega=(2,1,2,0,0,0,0,0,0,0)$). Illustration by the authors.
@@
Figure 4: a) A graphical representation of the cluster selection process in the case of coagulation, in which the probability of two particles merging does not depend on their sizes. Our task is to create a particle of size $g$, i.e., consisting of $g$ monomers. The monomers are distinguishable (numbered for clarity). In the first step, we combine two out of $g$ available particles (here $g=4$). In the next step, we combine two further freely selected clusters. We merge the particles until we obtain a single cluster. In this way, we implement one of $x_{4}=4!3!/2^{3}=18$ ways to create a cluster of size 44. b) Diagram illustrating all possible growth histories of such a particle of size 44. The number of different growth histories, $x_{g}$, equals the number of different paths drawn by arrows and leading through different states of the diagram. c) A tree corresponding to the bold path in the diagram in b). Single particle growth history may be illustrated as a rooted tree, with leaf nodes standing for monomeric units, internal nodes representing the history-dependent transition states of the cluster, and the root node being the final step in the growth process. Each vertical line represents a state of the system at a particular time step. Illustration by the authors.
II.4 Distributing monomers into clusters
@@
In this section, we will calculate the number of ways to distribute monomers into the given number of clusters (i.e., partitioning the system of NN monomers into $k$ subsets) Fronczak_2018 . As both $s$ and $g$ denote cluster size, at this point, we will use $g$, which will prove convenient in subsequent derivations, cf. Eqs. (22)-(26). Let NN distinguishable objects be divided into $k$ non-empty and disjoint subsets of $c_{i}>0$ elements each, where $\sum^{k}_{i=1}{c_{i}=N}$. As we can choose $c_{1}$ in $\binom{N}{c_{1}}$ ways, $c_{2}$ in $\binom{N-c_{1}}{c_{2}}$ ways, there are
@@
of such partitions; $n_{g}\geq 0$ stands for the number of subsets of size $g$, with the largest (possible) subset size equal to Nk+1N-k+1 (some of the subsets may stay empty). The change of the multiplication index $i$ to $g$ means that we no longer multiply over the subsets (i.e., over clusters) but over the sets of subsets (i.e., the sets of clusters) of a given size determined by $g$. For a better illustration of changing the index, see Fig. 3b.
For example, let us consider a system of $N=5$ monomers grouped into $k=2$ clusters of sizes 2 and 3. Remember that all monomers are distinguishable (i.e., we can assign numbers from 11 to $N=5$ to these particles). Possible grouping methods are presented in Fig. 3a. There are ten ways of doing so.
@@
${W_{p}(\mathrm{\Omega})=}N!\prod^{N-k+1}_{g=1}{\frac{1}{n_{g}!}{\left(\frac{x_{g}}{g!}\right)}^{n_{g}}}.$ (11)
Summing Eq. (11) over all integers $\left\{n_{g}\right\}$ specified by Eq. (4) one obtains the partial Bell polynomial $B_{N,k}\left(\left\{x_{g}\right\}\right)$, an important fact that will be used in Sec. II.7. Readers are encouraged to consult Appendix A for the definition and interpretation of partial Bell polynomials.
Internal state $x_{g}$ of the cluster, introduced in Eq. (11), in this context, simply corresponds to the number of ways in which the cluster of size $g$ could be created in the system (in other words, it is the number of possible histories of a given cluster). In general, $x_{g}$ varies for different kernels. The following subsection shows how to obtain $x_{g}$ for any given kernel.
To sum up this part, Eq. (11) gives the number of possible ways of partitioning NN monomers into $k$ clusters for a given state $\mathrm{\Omega}$ of the system. However, it implicitly assumes that all the clusters in the system arose at once, which is not valid, as only one merging act is allowed per time step. Coagulation acts corresponding to different clusters may alternate, increasing the number of ways in which a given (final) state can be created. This issue leads to the need to calculate the number of possible distributions of merging acts in time (see Sec. II.6).
II.5 Calculating number $x_{g}$ of ways to create cluster of given size for constant kernel
@@
where $x_{h}$ and $x_{g-h}$ represent the numbers of ways to create two clusters of size $h$ and (gh)(g-h) that joined and became a cluster of size $g$, and $x_{1}=1$.
@@
The second binomial factor, $\binom{g-2}{h-1}$, covers the fact that coagulation acts, resulting in clusters $h$ and (gh)(g-h), could occupy different possible time steps (cf. Fig. 5).
The sum in Eq. (13) is taken over all possible pairs of clusters that can join to become a cluster of size $g$. The factor of 1/2 is used to prevent double-counting of the coagulation acts. This factor is arbitrary in the sense of the final result, $\langle n_{s}\rangle$, which would not change if the factor of 1/2 were removed. More comments on this factor may be found in Ref. Lepek_2019 .
In the Appendix A.2, we show how to obtain Eq. (12) from Eq. (13) via the generating function method. Thus, we show that they define the same sequence. The recursive form will be later used to obtain $x_{g}$ for any given kernel.
II.6 Distributing joining acts in time
Figure 5: A graphical representation of the time evolution of a coalescing system (with a constant kernel) consisting of $N=7$ monomers, in which two particles with sizes 3 and 4 were formed. In the above figure, for illustration, we consider a state where the particle of size 3 was formed of monomers numbered as $\boxed{1}$, $\boxed{2}$, and $\boxed{3}$, and the particle of size 4 was formed of monomers $\boxed{4}$, $\boxed{5}$, $\boxed{6}$, and $\boxed{7}$ (monomers are distinguishable). For this selected case, which is one of the many possible microscopic realizations of the system, there are ten possible sequences of coagulation acts leading to the formation of these two clusters. To create them, we need five acts of coagulation, i.e., five time steps. The smaller cluster, size 3, is created in two time steps, and the larger cluster, size 4, is created in three time steps. In part a), we can see that the black dots, representing the coagulation acts that lead to the formation of a cluster of size 3, and the white dots, representing the coagulation acts that lead to the formation of a cluster of size 4, can be ordered in ten different ways in time. Part b) of the figure shows two of the ten possible system evolution scenarios. Coagulation acts associated with forming a cluster of size $h$ can occur in h1h-1 time steps out of the total number of N2=5N-2=5 steps needed to form clusters of sizes 3 and 4. For an arbitrary state $\Omega$, the total number of such sequences is given by Eq. (14), cf. Eq. (16). Reprinted from Lepek_2021 , with permission from Elsevier.
As signaled in the previous subsections, a given state can arise as a result of different sequences of intermixed coagulation acts corresponding to different clusters. In other words, if we study a given cluster, its related coagulation acts can occur in different steps in time. So far, we have not considered that fact. Therefore, the distribution of coagulation acts in time is the second origin of combinatorial equations.
@@ II.6 Distributing joining acts in time
$\begin{split}\left[\binom{t}{1}\binom{t-1}{1}\dots\binom{t-n_{2}+1}{1}\right]\dots\left[\binom{t-n_{2}}{2}\binom{t-n_{2}-2}{2}\dots\binom{t-n_{2}-2(n_{3}-1)}{2}\right]\dots{\\ }=\frac{t!}{{\left(1!\right)}^{n_{2}}{\left(2!\right)}^{n_{3}}\dots\left(\left(g-1\right)!\right)^{n_{g}}\dots}=t!\prod^{N-k+1}_{g=1}{\frac{1}{{\left(\left(g-1\right)!\right)}^{n_{g}}}}{=W_{t}(\mathrm{\Omega})}.\end{split}$ (14)
For a better understanding of Eq. (14), consider, as before, the system of $N=5$ monomers after $t=3$ time steps. Thus, $k=2$ clusters exist in that system. Let us assume those are: one cluster of size 2 (dimer) and one cluster of size 3 (trimer). From Eq. (14) we have
${W_{t}(\mathrm{\Omega})=W_{t}(\{0,1,1,0,0\})=}\frac{3!}{{\left(0!\right)}^{0}{\left(1!\right)}^{1}{\left(2!\right)}^{1}{\left(3!\right)}^{0}}=3,$ (15)
@@
$\begin{split}{W_{t}(\mathrm{\Omega})}&={W_{t}(\{0,0,1,1,0,0,0\})}\\ &=\frac{7!}{{\left(0!\right)}^{0}{\left(1!\right)}^{0}{\left(2!\right)}^{1}{\left(3!\right)}^{1}{\left(4!\right)}^{0}{\left(5!\right)}^{0}}={420}.\end{split}$ (16)
@@
Sequences that can be observed for a selected microscopic realization of that system are illustrated in Fig. 5.
It is worth noting that the form of Eq. (14) assumes that coagulation acts corresponding to different clusters may occupy unrestrictedly any of the available time steps; thus, it clearly results from the assumption of equiprobable ways of creation of a given state $\Omega$, Eq. (8).